OpenStudy (anonymous):

Find the inverse of the function: f(x) = 7x + 2

OpenStudy (anonymous):

Express x in terms of f(x) and numbers: 7x = f(x) - 2 x = [f(x0 - 2]/7

OpenStudy (slaaibak):

f-1(x) = x/7 -2/7

OpenStudy (anonymous):

also helps to think of it logically...if you take a number, multiply it by 7 and then add two, what would be the opposite process? taking a number, subtracting two, and then dividing by 7